OpenStudy (anonymous):

Please check my answer, A. @jdoe0001

OpenStudy (anonymous):

What do you think it is @lilia222 ??

OpenStudy (anonymous):

A. (As I state above ) lol.

OpenStudy (anonymous):

lol sry i feel dumb know

OpenStudy (anonymous):

Well good Job you are correct!!! :)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

how did you get A?

OpenStudy (anonymous):

One moment.

OpenStudy (jdoe0001):

\(\bf \textit{compounding continuously }=\large Pe^{rt}\to 54,200\cdot e^{0.072\cdot 21}\)

OpenStudy (anonymous):

I'm a bit confused now.

OpenStudy (anonymous):

i got A as well.

OpenStudy (jdoe0001):

the rate the restaurant attendance is increasing is a "continuously compounded" rate thus you'd use the "continously compound interest" formula

OpenStudy (jdoe0001):

\(\bf \textit{compounding continuously }=\large Pe^{rt}\to 54,200\cdot e^{0.072\cdot 21} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$54,200\\ r=rate\to 7.2\%\to \frac{7.2}{100}\to 0.072\\ t=years\to 21\)

OpenStudy (anonymous):

I got 245,839 rounded it would be B , correct?

OpenStudy (jdoe0001):

well... B is 58,239

OpenStudy (anonymous):

I meant C , my bad!!!

OpenStudy (jdoe0001):

but 245.839 is ... not really rounded up

OpenStudy (jdoe0001):

but yes, rounded up would be C

OpenStudy (anonymous):

245.839.9

OpenStudy (jdoe0001):

yeap.... which rounded is 245,840 :)

OpenStudy (anonymous):

So C is my final answer or do we have to do another step? O.o I swear that was it lol.

OpenStudy (jdoe0001):

yeap.. that's it

OpenStudy (anonymous):

:D Thank you!!! Do you mind if I tag you in more questions? **Are you going to be online? :/

OpenStudy (jdoe0001):

for a few, yes :)